OpenStudy (anonymous):

solve by the substitution method 3x-5y=-19 4x+31=y The solution is

OpenStudy (anonymous):

put y as 4x + 31 into the first equation 3x - 5(4x + 31) = -19 3x - 20 x -155 = -19 -17x = 136 x = -8

OpenStudy (anonymous):

y = 4x+31= 4(-8) + 31 = -1